Solve not in excel and provide all the formulas

You have been asked to evaluate the quality of investments at an some inc., a firm that open 2 business – books and pens. You have been provided the following information of the businesses:

Market value of assets

Unlevered beta of business

books

100 000 000$

1.30

pens

100 000 000$

0.70

200 000 000$

A firm has 100 million shares, trading at $10 a share and faces a marginal tax rate of 31% . Company`s bonds have an A rating and a dedefault spread ( over risk free rate) of 4%. The risk rate is 6% and the equity risk premium is 9%.

1) Market value of equity and market value of debt

2) Unlevered beta for the total firm ( can be found as a weighted average) and unlevered/levered beta for the firm

3) Cost of equity and debt

4) Cost of capital ( WACC)
Answer #1

1. The market value of equity = No. of shares y Trading price. here in the given case the No. of shares is given as 100 milli? Unleaveser beta of total from! Unlevered beta = a Beta of Books Market market value of Books & Mv od Books + MV od pen + -- 3 Cost of earity = Rg + ß (RM-R8.) Rm-2g = Risk premium = qy. Rg = Rist Ihoe rack = 6y. B = 1.69. = 6ot 1.69 x 9* = 21-21%05xloy Cr- 317) + 0.5 x 21:21% - 3.45 + 10.605 = 14. 0554.